# RoboMM: All-in-One Multimodal Large Model for Robotic Manipulation

### v1

@@ Abstract
In recent years, robotics has advanced significantly through the integration of larger models and large-scale datasets. However, challenges remain in applying these models to 3D spatial interactions and managing data collection costs. To address these issues, we propose the multimodal robotic manipulation model, RoboMM, along with the comprehensive dataset, RoboData. RoboMM enhances 3D perception through camera parameters and occupancy supervision. Building on OpenFlamingo, it incorporates Modality-Isolation-Mask and multimodal decoder blocks, improving modality fusion and fine-grained perception. RoboData offers the complete evaluation system by integrating several well-known datasets, achieving the first fusion of multi-view images, camera parameters, depth maps, and actions, and the space alignment facilitates comprehensive learning from diverse robotic datasets. Equipped with RoboData and the unified physical space, RoboMM is the generalist policy that enables simultaneous evaluation across all tasks within multiple datasets, rather than focusing on limited selection of data or tasks. Its design significantly enhances robotic manipulation performance, increasing the average sequence length on the CALVIN from 1.7 to 3.3 and ensuring cross-embodiment capabilities, achieving state-of-the-art results across multiple datasets.
Figure 1: Stacked bar chart depicting the performance of various models across five datasets. The “SOTA” label represents the best results achieved by specialized models for each dataset. Notably, RoboMM is the only generalist policy evaluated on multiple datasets, demonstrating competitive performance relative to “SOTA”. For more details, please refer to Section 5.1.
@@ 1 Introduction
In recent years, machine learning has experienced profound advancements, from the advent of CLIP [53, 67] to the progression of foundational models like the GPT series [8, 1, 9], Llama [59, 60], LLaVA [37], and Flamingo [2, 3]. These strides are largely due to larger transformer-based architectures and the utilization of “internet-scale” datasets [36, 13, 11, 14]. These innovations have not only extended the frontiers of natural language processing [16] and computer vision [22, 54] but have also galvanized researchers to integrate these models into Embodied Artificial Intelligence (EAI) [49], thereby enabling more complex and varied tasks in real-world environments.
In terms of modeling, there has been a gradual shift from single-task or single-dataset learning [57, 42, 67] towards transfer learning approaches [7, 30, 35, 61, 58, 25]. These models leverage robust foundational models that are pre-trained on extensive “internet-scale” datasets or diverse data sources. Subsequently, they undergo fine-tuning on specific robotic datasets to produce precise control actions. On the data front, researchers has collected data through various means to augment models. For example, the Open X-Embodiment [51] amalgamates diverse datasets containing vision-language-action pairs, whereas the RH20T [18] gathers data via teleoperation. Despite the impressive robustness of these efforts, they still encounter significant challenges in practical applications.
@@
Figure 2: Left: RoboData integrates nine diverse and complex datasets (CALVIN [41], Meta-World [65], LIBERO [34], Robomimic [39], RoboCAS [68], ManiSkill2 [21], RoboCasa [47], RLBench [26], and Colosseum [52]), covering various robot embodiments, environments, and task types. Standardized input and output spaces achieve a unified dataset. Right: RoboMM features comprehensive 3D environmental perception capabilities, flexible and diverse multimodal outputs, and significantly enhances the robotic manipulation generalization capabilities.
Firstly,  is the direct application of multimodal models to EAI the optimal solution? Robots must interact with the physical 3D space; however, current multimodal models predominantly focus on 2D image understanding and generation, which may limit their practical applicability. Secondly,  is it essential to address the cost and efficiency of dataset construction? For instance, collecting around 130,000 episodes from the RT-1 [6] took 17 months. Therefore, it is imperative to integrate as many existing multi-platform, multi-robot datasets from the industry as possible to address this urgent issue.
@@
In response to these challenges, this paper introduces RoboMM, the native multimodal large model for robotic manipulation, and RoboData, the comprehensive dataset integrating datasets across various platforms and robots for evaluation and training purposes.
@@
In terms of modeling, RoboMM combines camera parameters and occupancy supervision to enhance 3D environmental perception. Additionally, leveraging large language models like OpenFlamingo [3], we design the plug-and-play efficient Modality-Isolation-Mask (MIM), which flexibly introduces multimodal supervision. This not only grants the model fine-grained perception capabilities but also enables the effective utilization of vast amounts of internet data. On the data side, while Open X-Embodiment [51] integrates multiple datasets, it lacks critical information such as multi-view images, camera parameters, and depth maps, making it more suitable for 2D multimodal training. Additionally, the lack of data space alignment prevents the robot’s 6D pose from being consistent across different datasets. In contrast, RoboData addresses these limitations by integrating a wide range of well-known industry datasets, including CALVIN [41], Meta-World [65], LIBERO [34], Robomimic [39], RoboCAS [68], ManiSkill2 [21], RoboCasa [47], RLBench [26], and Colosseum [52], which are not covered in Open X-Embodiment [51]. We dedicate 200 person-days to collect and organize these data, supplementing missing modalities such as depth maps and camera parameters. Furthermore, thanks to the unified physical space, RoboData aligns the input and output spaces of cross robots and platforms, ensuring consistency and facilitating integrated learning from diverse robotic datasets.
RoboData aims to provide the industry with a comprehensive and fair evaluation system, while RoboMM is the generalist policy to incorporate training and testing across multiple datasets. Extensive experiments demonstrate that the various components of RoboMM significantly improve performance in robotic manipulation tasks, enhancing the average sequence length on the CALVIN [41] benchmark from 1.7 to 3.3. Additionally, RoboMM ensures cross-embodiment capabilities and achieves state-of-the-art results across multiple datasets, as shown in Figure 1. This paper underscores the critical role of advanced models and curated datasets in advancing robotics, highlighting their potential to drive significant improvements in robotic performance and functionality.
@@
Figure 3: Architecture of RoboMM. Vision Encoder Block for extracting multi-view features, Adapter Block that leverages occupancy supervision to unify features and enhance spatial perception, Feature Fusion Block based on LLMs for merging text and visual information and Multimodal Decoder Blocks that enhance fine-grained perception and understanding through multimodal outputs.
2 Related Work
3 RoboMM
3.1 Review
@@
3.2 Network Architecture
For robotic manipulation tasks based on language instructions $T$, which typically rely on historical frames $I$ from $N$ perspectives at $H$ time steps, the task can be mathematically expressed as $O_{A}=\Theta(I,T)$, where $I\in\mathbb{R}^{H\times N\times H\times W\times 3}$. Integrating the principles of MLLMs, this paper proposes the novel native multimodal robotic manipulation model named RoboMM, as show Figure 3, which has the capability of 3D environment perception and handles multimodal inputs (text $T$, vision $I$, camera parameters $Cam$) and outputs (actions $O_{A}$, images $O_{I}$, occupancy $O_{o}$):
$(O_{A},[O_{I},O_{O}])=\text{RoboMM}(T,I,Cam).$ (2)
RoboMM consists of the following key components: (1) Vision Encoder Block: Designed to extract observation features $F_{I}^{h,n}$ from $H$ time steps and $N$ perspectives. (2) 3D Perception Adapter Block: Enhances physical space perception by integrating camera parameters. (3) Feature Fusion Decoder Based on Large Language Models: Merges text and visual information to output multimodal features, with the use of Modality-Isolation-Mask (MIM) increasing the flexibility of modality fusion. (4) Multimodal Decoder Blocks: Enhances the model’s fine-grained perception and understanding through multimodal outputs. Notably, thanks to MIM, $O_{I},O_{o}$ are optional outputs.
Adapter: We employ UVFormer from RoboUniView [35], which is a simple yet powerful 3D environment perception model. UVFormer takes image features $X^{h}=\{F_{I}^{h,n}\}_{n}^{N}$, camera parameters $Cam^{h}=\{Cam^{h,n}\}_{n}^{N}$, and learnable unified view queries $Q$ as inputs and outputs the unified view representation $U_{I}^{h}$:
@@ 3.2 Network Architecture
Feature Fusion Decoder: Due to the need to support multi-frame or video inputs, we abandon the Auto-Regressive (AR) mechanism used in LLaVA [36] and adopt OpenFlamingo [3] with cross-attention as the Feature Fusion Decoder. It integrates unified visual representations with language and other modality placeholders through cross-attention layers.
@@
[…]
@@
(b) Attention Fusion: Continuing with the use of cross-attention in OpenFlamingo [3], we fuse visual and text features wherein the text features $F^{h}_{T}$ serve as the query, and the visual features $U_{I}^{h}$ serve as the key and value. It is worth noting that the self-attention layer incorporates MIM(see in Figure 5 Left), which allows training with auxiliary modality supervision and omitting unnecessary modalities during inference, significantly increasing the flexibility of modality fusion.
@@
(a) Image Decoder: As shown in Figure 4(a), we design a simple structure that includes 2 attention decoder layers. This structure outputs image patches, which are then assembled into a complete image(static images $O^{h}_{simg}$ or wrist image $O^{h}_{gimg}$) based on their coordinates.
(b) Occupancy Decoder: The initial part of this structure, as shown in Figure 4(b), is similar to Image Decoder, generating the feature $U^{h}_{occ}$. Then, $U^{h}_{occ}$ is reshaped, upsampled, and processed through 3D convolutions to reconstruct the entire 3D occupancy $O^{h}_{o}=\{o^{h}_{pos},o^{h}_{rgb}\}$. The flexible model architecture allows the visual module to generate the occupancy map $O^{h}_{o}=O^{h}_{o_{v}}$ from multi-view features using UVFormer, while LLM also outputs $O^{h}_{o}=O^{h}_{o_{m}}$, corresponding to $T_{occ}$. Experimental validation shows that $O^{h}_{o_{v}}$ and $O^{h}_{o_{m}}$ provide similar assistance for robotic manipulation. Unless otherwise specified, in this paper, $O^{h}_{o}=O^{h}_{o_{v}}$.
(c) Action Decoder: In Figure 4(c), we use a few MLP layers to output actions $O^{h}_{A}$ consisting of delta 6D poses $a_{pose}^{h}$ = $\Delta\text{pos}_{x}^{h},\Delta\text{pos}_{y}^{h},\Delta\text{pos}_{z}^{h},\Delta\text{rot}_{x}^{h},\Delta\text{rot}_{y}^{h},\Delta\text{rot}_{z}^{h}\}$ and 1-DoF gripper actions $a_{g}^{h}$.
@@ 3.3 Training Objective
Here, $\lambda_{\text{image}}$ and $\lambda_{\text{occ}}$ are weight coefficients for image and occupancy losses. Notably, $l_{simg}$, $l_{gimg}$, and $l_{o}$ can be excluded from training if the corresponding modality is unavailable, providing a flexible optimization framework for multimodal learning.
Action Loss $l_{a}$: The action loss function optimizes $a_{pose}^{h}$ and $a_{gripper}^{h}$ using the combination of Mean Squared Error (MSE) and Binary Cross-Entropy (BCE):
@@
Image Loss $l_{simg}$ or $l_{gimg}$: The image loss measures pixel-level differences between the predicted images and the next frame images $\hat{I}^{h+1}_{simg}$ or $\hat{I}^{h+1}_{gimg}$ using L2 loss:
@@
This effectively guides the model in predicting future scenes.
@@ 4 RoboData
To address these challenges, we curate well-known datasets from the industry, including CALVIN [41], Meta-World [65], LIBERO [34], Robomimic [39], RoboCasa [47], ManiSkill2 [21], RoboCAS [68], RLBench [26], and Colosseum [52], forming a comprehensive dataset we call RoboData. This dataset aims to provide the industry with a complete and fair evaluation system, comprising 70,000 episodes and 7 million samples. It encompasses a diverse range of tasks, including placing, picking, turning, and stacking. Figure 5 Right illustrates these tasks along with their corresponding number of episodes, highlighting the distribution of tasks within the dataset.
@@
3D Space Alignment: We focus on the unification of world coordinate systems, workspaces, and action spaces. Different datasets adopt their own coordinate systems. For example, the RLBench [26] and Colosseum [52] reference the robot’s body, setting the X-axis to point forward, the Y-axis to point left, and the Z-axis to point upward; whereas ManiSkill2 [21] orient the X-axis forward, the Y-axis right, and the Z-axis downward. Figure 6 illustrate the movements of robotic arms from RLBench [26] and ManiSkill2 [21] datasets, respectively. Although both exhibit similar motion directions (moving from top to bottom), the representation of actions differs significantly due to the variations in coordinate systems. For instance, in RLBench [26], $a_{pose}=[0.0,0.0,-0.1,0.0,0.0,0.0]$, while in ManiSkill2 [21], $a_{pose}=[0.0,0.0,0.1,0.0,0.0,0.0]$.
@@
Action Representation Alignment: Various datasets employ different methods to obtain robot actions, and this diversity can lead to inconsistencies in the data. For instance, CALVIN [41] uses the Euler Angle Difference Method (EADM) for action representation, while LIBERO [34], Robomimic [39], RoboCasa [47] utilize the Composite Rotation Matrix Method (CRMM), and ManiSkill2 [21] adopts the Pose Composition Method (PCM). To address this issue, we unify the action representation across all datasets by regenerating them using the CRMM, which has been validated in SRT [29]. This choice not only enhances data consistency but also provides a more reliable foundation for subsequent research. For more detailed technical information, please refer to Chapter 8.
Missing Data Imputation: In many datasets, camera intrinsics and extrinsics are often not directly provided, which poses challenges for research. To tackle this problem, we reconstruct the original simulations and render them anew, leveraging the original data to drive the robots in order to obtain these critical parameters. This process ensures data completeness and provides essential support for subsequent experiments.
In conclusion, by effectively aligning the input and output spaces, we propose a new dataset standard aimed at comprehensively optimizing existing datasets. The implementation of this standard will facilitate the development of more versatile and general-purpose embodied AI agents. We plan to release this dataset along with evaluation code to promote further research and advancement in the field of embodied intelligence. Additionally, we welcome all like-minded researchers to join our efforts in driving the development of this field.
@@
CALVIN MDT [55] RSS’24 93.7%
HULC++ [43] ICRA’24 93.0%
SPIL [70] RA-L’24 84.6%
LCD [66] arXiv’23 88.7%
RoboFlamingo [33] arXiv’23 86.0%
PlayFusion [12] CoRL’23 45.2%
Distill-D [24] CoRL’23 86.7%
HULC [40] RA-L’22 82.7%
CALVIN [41] RA-L’22 76.4%
RoboMM (ours) - 91.0%
RoboMM- (ours) - 74.7%
Meta-World PRISE [69] ICML’24 80.4%
PAD [23] NeurIPS’24 72.5%
GR-1 [64] ICLR’24 57.4%
SuSIE [5] ICLR’24 41.0%
RT-2* [7] arXiv’23 52.2%
RT-1 [6] RSS’23 34.6%
RoboMM (ours) - 78.6%
RoboMM- (ours) - 79.3%
@@
RoboMM (ours) - 90.7%
RoboMM- (ours) - 64.2%
@@
RoboMM (ours) - 30.6%
RoboMM- (ours) - 27.0%
Robomimic IBC [19] CoRL’21 13.6%
RoboMM (ours) - 15.0%
RoboMM- (ours) - 8.0%
Table 1: Performance on Various Datasets.

### v2

# RoboTron-Mani: All-in-One Multimodal Large Model for Robotic Manipulation

@@ Abstract
Recently, robotics has advanced significantly through the integration of larger models and large-scale datasets. However, challenges remain in applying these models to 3D spatial interactions and managing data collection costs. To address these issues, we propose the multimodal robotic manipulation model RoboTron-Mani and the comprehensive dataset RoboData. RoboTron-Mani, on one hand, enhances 3D perception through camera parameters and occupancy supervision. On the other hand, it further incorporates Modality-Isolation-Mask and multimodal decoder blocks based on OpenFlamingo, improving modality fusion and fine-grained perception. RoboData integrats several publicly-available datasets, achieving the first fusion of multi-view images, camera parameters, depth maps, actions, and space alignment, which facilitates comprehensive learning from diverse robotic datasets and offers one complete evaluation system. Trained on RoboData, RoboTron-Mani is the first generalist policy that surpasses expert models, enabling simultaneous evaluation of all tasks across multiple datasets, rather than being limited to specific data or task selections. Specifically, RoboTron-Mani boosts manipulation performance by increasing the average sequence length on CALVIN from 1.7 to 3.5, enabling cross-embodiment generalization, and achieving state-of-the-art results on both simulated and real-world datasets.
Figure 1: Stacked bar chart depicting the performance of various models across five datasets. The “SOTA” label represents the best results achieved by specialized models for each dataset. Notably, RoboTron-Mani is the only generalist policy evaluated on multiple datasets, demonstrating competitive performance relative to “SOTA”. For more details, please refer to Section 4.1.
@@ 1 Introduction
In recent years, machine learning has experienced profound advancements, from the advent of CLIP [56, 70] to the progression of foundational models like the GPT series [10, 1, 11], Llama [62, 63], LLaVA [39], and Flamingo [2, 3]. These strides are largely due to larger transformer-based architectures and the utilization of “internet-scale” datasets [38, 15, 13]. These innovations have not only extended the frontiers of natural language processing [17] and computer vision [23, 57] but also galvanized researchers to integrate these models into Embodied Artificial Intelligence (EAI) [51], thereby enabling more complex and varied tasks in real-world environments.
@@
Figure 2: Left: RoboData integrates multiple diverse and complex datasets (CALVIN [43], Meta-World [68], LIBERO [36], RT-1 [8], RoboCAS [71], ManiSkill2 [22], RoboCasa [49], RLBench [27], and Colosseum [55]), covering various robot embodiments, environments, and task types, resulting in a unified dataset with standardized input and output spaces. Right: RoboTron-Mani features comprehensive 3D perception capabilities, flexible multimodal outputs, and significantly enhances the robotic manipulation generalization capabilities.
In terms of modeling, there has been a gradual shift from single-task or single-dataset learning [60, 44, 70] towards transfer learning approaches [9, 31, 37, 64, 61, 26]. These models leverage robust foundational models that are pre-trained on extensive “internet-scale” datasets or diverse data sources. Subsequently, they undergo fine-tuning on specific robotic datasets to produce precise control actions. On the data front, researchers collect data through various means to augment models. For example, the Open X-Embodiment [53] amalgamates diverse datasets containing vision-language-action pairs, whereas the RH20T [20] gathers data via teleoperation. Despite the impressive robustness of these efforts, they still encounter significant challenges in practical applications. First,  is the direct application of multimodal models to EAI the optimal solution? Robots should interact with the physical 3D space. However, current multimodal models predominantly focus on 2D image understanding and generation, limiting their practical applicability, especially in physical-world interaction. Second,  is it essential to address the cost and efficiency of dataset construction? For instance, collecting around 130,000 episodes from the RT-1 [8] took 17 months. Therefore, it is imperative to integrate as many existing multi-platform, multi-robot datasets from the industry as possible to address this urgent issue.
@@
For tackling the aforementioned challenges, this paper introduces RoboTron-Mani, the multimodal large model for robotic manipulation, and RoboData, the comprehensive dataset integrating datasets across various platforms and robots for evaluation and training purposes. Figure 2 shows the framework and capabilities of our dataset platform and multimodal robot model.
In terms of modeling, RoboTron-Mani combines camera parameters and occupancy supervision to enhance 3D environmental perception. Additionally, leveraging large language models like OpenFlamingo [3], we design the plug-and-play efficient Modality-Isolation-Mask (MIM), which flexibly introduces multimodal supervision. This gives the model fine-grained perception and leverages large-scale internet data.
@@
On the data side, while Open X-Embodiment [53] integrates multiple datasets, it lacks critical information such as multi-view images, camera parameters, and depth maps, making it more suitable for 2D multimodal training. Additionally, the lack of data space alignment prevents the robot’s 6D pose from being consistent across different datasets, yielding the results in Open X-Embodiment [53] Table 1, where direct data fusion results in RT-1-X underperforming RT-1. In contrast, RoboData addresses these limitations by integrating a wide range of well-known industry datasets, including simulated datasets (CALVIN [43], Meta-World [68], LIBERO [36], Robomimic [41], RoboCAS [71], ManiSkill2 [22], RoboCasa [49], RLBench [27], and Colosseum [55]), and RT-1 [8]. We dedicate hundrads of person-days to collect and organize these data, supplementing missing modalities such as depth maps and camera parameters. Moreover, RoboData unifies the input and output spaces across different robots and platforms, ensuring consistent perception and action representation, and enabling integrated learning from diverse simulated and real-world datasets.
@@
Figure 3: Architecture of RoboTron-Mani. Vision Encoder extracts multi-view features. 3D Perception Adapter leverages occupancy supervision to unify features and enhance spatial perception. Feature Fusion Decoder based on LLMs merges text and visual information. Multimodal Decoders enhance fine-grained perception and understanding through multimodal outputs.
Extensive experiments demonstrate that the various components of RoboTron-Mani significantly improve performance in robotic manipulation tasks, enhancing the average sequence length on the CALVIN [43] benchmark from 1.7 to 3.5. Additionally, RoboTron-Mani ensures cross-embodiment capabilities and the first generalist policy that surpasses expert models, enabling simultaneous evaluation of all tasks across multiple datasets, rather than being limited to specific data or task selections, as shown in Figure 1. This paper underscores the critical role of advanced models and curated datasets in advancing robotics, highlighting their potential to drive significant improvements in robotic performance. In summary, RoboData aims to provide the research community with a comprehensive and fair evaluation system. And RoboTron-Mani is the first generalist policy to incorporate training and testing across multiple datasets.
2 RoboTron-Mani
2.1 Preliminary
@@
2.2 Architecture
For robotic manipulation tasks based on language instructions $T$, which typically rely on historical frames $I$ from $N$ perspectives at $H$ time steps, the task can be mathematically expressed as $O_{A}=\Theta(I,T)$, where $I\in\mathbb{R}^{H\times N\times H\times W\times 3}$. Integrating the principles of MLLMs, this paper proposes the novel multimodal robotic manipulation model named RoboTron-Mani, as show Figure 3, which has the capability of 3D environment perception and handles multimodal inputs (text $T$, vision $I$, camera parameters $Cam$) and outputs (actions $O_{A}$, images $O_{I}$, occupancy $O_{o}$):
$(O_{A},[O_{I},O_{O}])=\text{RoboTron-Mani}(T,I,Cam).$ (2)
RoboTron-Mani consists of the following key components: (1) Vision Encoder extracts observation features $F_{I}^{h,n}$ from $H$ time steps and $N$ perspectives. (2) The 3D Perception Adapter enhances spatial understanding by incorporating camera parameters.(3) Feature Fusion Decoder Based on Large Language Models merges text and visual information to output multimodal features, with the use of Modality-Isolation-Mask (MIM) increasing the flexibility of modality fusion. (4) Multimodal Decoder enhances the model’s fine-grained perception and understanding through multimodal outputs. Notably, thanks to MIM, $O_{I},O_{o}$ are optional outputs.
3D Perception Adapter. We employ UVFormer from RoboUniView [37], which is a simple yet powerful 3D environment perception model. UVFormer takes image features $X^{h}=\{F_{I}^{h,n}\}_{n}^{N}$, camera parameters $Cam^{h}=\{Cam^{h,n}\}_{n}^{N}$, and learnable unified view queries $Q$ as inputs and outputs the unified view representation $U_{I}^{h}$:
@@ 2.2 Architecture
Feature Fusion Decoder. Due to the need to support multi-frame or video inputs, we abandon the Auto-Regressive (AR) mechanism used in LLaVA [38] and adopt OpenFlamingo [3] with cross-attention as the Feature Fusion Decoder. It integrates unified visual representations with language and other modality placeholders through cross-attention layers.
@@
$T^{\prime}=\{$ $[T_{\text{img}},T^{h},T_{\text{simg}},T_{\text{gimg}},T_{\text{occ}},T_{\text{act}}]\}_{h}^{H}.$ (4)
@@
(b) Attention Fusion: Continuing with the use of cross-attention in OpenFlamingo [3], we fuse visual and text features wherein the text features $F^{h}_{T}$ serve as the query, and the visual features $U_{I}^{h}$ serve as the key and value. It is worth noting that the self-attention layer incorporates MIM (Figure 5 Left), which allows training with auxiliary modality supervision and omitting unnecessary modalities during inference, significantly increasing the flexibility of modality.
@@
(a) Image Decoder. As shown in Figure 4(a), we design a simple structure that includes 2 attention decoder layers. This structure outputs image patches, which are then assembled into a complete image (static images $O^{h}_{simg}$ or wrist image $O^{h}_{gimg}$) based on their coordinates.
(b) Occupancy Decoder. The initial part of this structure, as shown in Figure 4(b), is similar to Image Decoder, generating the feature $U^{h}_{occ}$. Then, $U^{h}_{occ}$ is reshaped, upsampled, and processed through 3D convolutions to reconstruct the entire 3D occupancy $O^{h}_{o}=\{o^{h}_{pos},o^{h}_{rgb}\}$. The flexible model architecture allows the visual module to generate the occupancy map $O^{h}_{o}=O^{h}_{o_{v}}$ from multi-view features using UVFormer, while LLM also outputs $O^{h}_{o}=O^{h}_{o_{m}}$, corresponding to $T_{occ}$. Experimental validation shows that $O^{h}_{o_{v}}$ and $O^{h}_{o_{m}}$ provide similar assistance for robotic manipulation. Unless otherwise specified, in this paper, $O^{h}_{o}=O^{h}_{o_{v}}$.
(c) Action Decoder. In Figure 4(c), we use a few MLP layers or DiT [34] blocks to output actions $O^{h}_{A}$, consisting of delta 6D poses $a_{pose}^{h}$ = $\Delta\text{pos}_{x}^{h},\Delta\text{pos}_{y}^{h},\Delta\text{pos}_{z}^{h},\Delta\text{rot}_{x}^{h},\Delta\text{rot}_{y}^{h},\Delta\text{rot}_{z}^{h}\}$ and 1-DoF gripper actions $a_{g}^{h}$. Unless specified otherwise, this paper defaults to using an MLP action head. For more details on DiT, please refer to CogACT [34].
@@ 2.3 Training Objective
Here, $\lambda_{\text{image}}$ and $\lambda_{\text{occ}}$ are balancing weights for image and occupancy losses. Notably, $l_{simg}$, $l_{gimg}$, and $l_{o}$ can be excluded from training if the corresponding modality is unavailable, providing a flexible optimization framework.
Action Loss $l_{a}$. The action loss function optimizes $a_{pose}^{h}$ and $a_{gripper}^{h}$ using the combination of Mean Squared Error (MSE) and Binary Cross-Entropy (BCE):
@@
Image Loss $l_{simg}$ or $l_{gimg}$. The image loss measures pixel-level differences using L2 loss between the predicted images and the next frame images $\hat{I}^{h+1}_{simg}$ or $\hat{I}^{h+1}_{gimg}$:
@@ 3 RoboData
To address these challenges, we curate well-known publicly available datasets, including CALVIN [43], Meta-World [68], LIBERO [36], Robomimic [41], RoboCasa [49], ManiSkill2 [22], RoboCAS [71], RLBench [27], Colosseum [55] and RT-1 [8], forming a comprehensive dataset we call RoboData, aiming to provide the industry with a complete and fair evaluation system. Specifically, RoboData comprises 70,000 episodes and 7 million samples, and encompasses a diverse range of tasks, including placing, picking, turning, and stacking. Figure 5 Right illustrates these tasks along with their corresponding number of episodes, highlighting the distribution of tasks within the dataset.
@@
3D Space Alignment. We focus on the unification of world coordinate systems, workspaces, and action spaces. Different datasets adopt their own coordinate systems. For example, the RLBench [27] and Colosseum [55] reference the robot’s body, setting the X-axis to point forward, the Y-axis to point left, and the Z-axis to point upward; whereas ManiSkill2 [22] orient the X-axis forward, the Y-axis right, and the Z-axis downward. Figure 6 illustrate the movements of robotic arms from RLBench [27] and ManiSkill2 [22] datasets, respectively. Although both exhibit similar motion directions (moving from top to bottom), the representation of actions differs significantly due to the variations in coordinate systems. For instance, in RLBench [27], $a_{pose}=[0.0,0.0,-0.1,0.0,0.0,0.0]$, while in ManiSkill2 [22], $a_{pose}=[0.0,0.0,0.1,0.0,0.0,0.0]$.
@@
Action Representation Alignment. Various datasets employ different methods to obtain robot actions, and such diversities lead to inconsistencies in the data. For instance, CALVIN [43] uses the Euler Angle Difference Method (EADM) for action representation, while LIBERO [36], Robomimic [41], RoboCasa [49] utilize the Composite Rotation Matrix Method (CRMM), and ManiSkill2 [22] adopts the Pose Composition Method (PCM). To address this issue, we unify the action representation across all datasets by regenerating them using the CRMM, which has been validated in SRT [30]. This choice not only enhances data consistency but also provides a more reliable foundation for subsequent research. For more detailed technical information, please refer to Chapter 9.
Missing Data Imputation. In many datasets, camera intrinsics and extrinsics are often not directly provided, which poses challenges for research. To tackle this problem, we reconstruct the original simulations and render them anew, leveraging original data to drive the robots in order to obtain these critical parameters. This process ensures data completeness and provides essential support for experiments.
In summary, by effectively aligning the input and output spaces, we propose a new dataset standard aimed at comprehensively optimizing existing datasets. The implementation of this standard will facilitate the development of more versatile and general-purpose embodied AI agents. We plan to release this dataset along with evaluation code to promote further research and advancement in the field of embodied intelligence. Additionally, we welcome all researchers to join our efforts in driving the development of this field.
@@
RoboFlamingo [35] arXiv’23 72.1%
$\pi 0^{*}$ [6] arXiv’24 87.3%
$\pi$-Fast [54] arXiv’25 89.1%
RoboTron-Mani(ours) - 91.7%
@@
RoboFlamingo [35] arXiv’23 15.6%
GR00T-N1-2B+IDM [5] arXiv’25 40.9%
RoboTron-Mani(ours) - 47.4%
CALVIN DTP [19] arXiv’25 92.4%
MDT [58] RSS’24 93.7%
HULC++ [45] ICRA’24 93.0%
SPIL [73] RA-L’24 84.6%
LCD [69] arXiv’23 88.7%
RoboFlamingo [35] arXiv’23 86.0%
PlayFusion [14] CoRL’23 45.2%
Distill-D [25] CoRL’23 86.7%
HULC [42] RA-L’22 82.7%
CALVIN [43] RA-L’22 76.4%
RoboFlamingo [35] arXiv’23 43.8%
RoboTron-Mani(ours) - 93.8%
Meta-World PRISE [72] ICML’24 80.4%
PAD [24] NeurIPS’24 72.5%
GR-1 [67] ICLR’24 57.4%
SuSIE [7] ICLR’24 41.0%
RT-2* [9] arXiv’23 52.2%
RT-1 [8] RSS’23 34.6%
RoboFlamingo [35] arXiv’23 65.3%
RoboTron-Mani(ours) - 80.1%
RT-1 Dataset RT-2-X(55B) [53] ICRA’24 60.7%
RT-1-X [53] ICRA’24 53.4%
Octo-Base [61] RSS’24 16.9%
OpenVLA [31] arXiv’24 24.8%
HPT [64] NeurIPS’24 35.0%
RoboTron-Mani(ours) - 60.0%
Table 1: Performance on Various Datasets. Bold indicates the best result, underline indicates the second-best result, and denotes reproduced results, where RoboFlamingo is the result of joint training on four simulation datasets.